Half life of a radioactive substance is 33.6 years
a)find rate of decay
b) how much of the sample of 400 grams will remain after 20 years?
c)how long will it take a sample of 400 grams to decay to 300?

Answer by stanbon(75887)

Half life of a radioactive substance is 33.6 years
----
1/2Ao = Ao(1/2)^(t/33.6)
------
a) find rate of decay
rate = 1 - (1/2)^(1/33.6) = 1-0.9796 = -0.0204
--------
b) how much of the sample of 400 grams will remain after 20 years?
x = 400(1/2)^(20/33.6)
x = 254.77 grams
------------------------
c)how long will it take a sample of 400 grams to decay to 300?
300 = 400(1/2)^(t/33.6)
(1/2)^(t/33.6) = 0.75
(t/33.6)log(0.5) = log(0.75)
---
t/33.6 = 0.4150
----
t = 13.95 years
========================
Cheers,
Stan H.
==================